A car is parked on a cliff overlooking the ocean on an incline that makes an angle of 24 degrees below the horizontal. The car rolls from rest down the incline with a constant acceleration of 4 ms^-2 for a distance of 50 m to the edge of the cliff, which is 30 m above the ocean. Find:
a) the car’s position relative to the base of the cliff when the car lands in the ocean
b) the length of time the car is in the air
1
Expert's answer
2018-08-06T10:52:08-0400
s=(at^2)/2
t=√(2(50)/4)=5 s
v=at=(4)5=20 m/s
h(0)=30
h(t)=-(9.8t^2)/2-20 sin⁡24 t+30=0
t=1.78 s
d=20 cos⁡24 (1.78)=32.5 m.
b)
t=1.78 s